OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means. WILL FAN AND GIVE MEDAL

OpenStudy (anonymous):

PLS HELP.

OpenStudy (anonymous):

-3 is the common difference, but I just need help explaining the formula of finding that out.

OpenStudy (anonymous):

given sequence = 4 + 1 +(-2) + (-5) +(-8) +(-11) thus this means a0 = 4 a1 = 1 a2 = -2 and so on as it is given in question that this series is an arithmetic series thus a1 = a0+d 1 = 4 + d => d = -3 a2 = a0+ 2d -2 = 4 + 2d -6 = 2d -3 = d now use the formula of ap series sum and find the sum

OpenStudy (anonymous):

Thank you so much! :)